# Variable Elliptic Structures on the Plane: Transport Dynamics, Rigidity, and Function Theory

### v1

@@ Introduction
The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work provides concrete examples of this conservative regime. These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems.
@@ 1.3 Differentiating the generator
A first concern in any variable-coefficient theory is whether the moving generator $i=i(x,y)$ can be differentiated in a meaningful and computable way. In our setting $i$ is a $C^{1}$ section of the algebra bundle, hence its partial derivatives $i_{x},i_{y}$ are well-defined as sections as well. In the elliptic regime, these derivatives can be isolated explicitly from the structure reduction. More precisely, the computation relies only on the invertibility of the element $2i+\beta$ in the underlying algebra. This condition is satisfied not only in the elliptic case but also in the hyperbolic and parabolic regimes, away from zero divisors. Consequently, the algebraic resolution of $i_{x}$ and $i_{y}$ from the structure reduction can be performed verbatim in all three cases; what distinguishes the elliptic regime is not the computation itself but the positivity and invertibility properties that follow from it.
@@ 1.6 Explicit Formula for the Inhomogeneous Coefficients
Differentiating the structure reduction (1.3) gives
@@ 2.3 Eliminating the Derivatives of the Generator
This system is valid for every variable elliptic structure. No compatibility or integrability assumption has been used.
@@ 2.4 Canonical Spectral Parameter
This change of variables depends only on the structure reduction. It introduces no new assumptions and does not involve differentiation.
@@ 3.2 The Obstruction Field
and that its derivatives $i_{x}$ and $i_{y}$ are well-defined algebra-valued sections.
@@ 3.4 Rigidity and Conservative Transport
$\lambda_{x}+\lambda\,\lambda_{y}=\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1},$
where the forcing terms $(g_{0},g_{1})$ are the real coefficients of where $\mathcal{G}=G_{0}+G_{1}i$ is the decomposition of the obstruction field in the moving basis as in (2.2).
@@ 3.6 Rigidity Theorem
admits an equivalent intrinsic reformulation. It states that the generator $i(x,y)$, viewed as a purely imaginary section of the rank–two algebra bundle it generates, is holomorphic with respect to its own variable elliptic structure:
@@ 3.7 Interpretation
This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores integrability within an already dynamical framework.
In the next chapter we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
@@ 4.1 Purpose and Scope
Chapter 2 established that variable elliptic structures are governed, at the structural level, by a transport equation for their spectral data. Chapter 3 identified rigidity as the distinguished regime in which this transport becomes conservative.
The purpose of the present chapter is twofold:
  • to analyze the conservative transport law as a first–order dynamical system;
  • to construct explicit, genuinely nonconstant rigid elliptic structures.
No new structural assumptions are introduced. All results here are consequences of the conservative Burgers equation and are independent of analytic considerations such as integral formulas or function spaces.
@@ 4.3 Characteristic Formulation
Equivalently, along curves of slope $\lambda$, the spectral parameter is conserved. This immediately implies:
@@ 4.4 Real Coefficient Representation
This system governs the evolution of the structure coefficients $(\alpha,\beta)$ in the rigid regime. It is quasilinear, first order, and hyperbolic.
@@ 4.6 Reduction to Ordinary Differential Equations
Then system (4.3) reduces to
$a_{x}=0,\qquad b_{x}=0,$
along characteristics unless $a$ depends on $x$ alone. The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
@@ 4.8 The $\varepsilon$–Family of Rigid Structures
For every $\varepsilon\in\mathbb{R}$, the coefficients (4.6) define a rigid elliptic structure on the domain $\{1-\varepsilon x\neq 0\}$.

### v2

@@ Introduction
The explicit $\varepsilon$–family of rigid elliptic structures constructed in this work by standard techniques provides concrete examples of this conservative regime. These examples show that rigidity does not force triviality: genuinely nonconstant structures exist, although ellipticity naturally restricts the domain of definition. From the transport viewpoint, these restrictions reflect characteristic phenomena familiar from first–order hyperbolic systems.
@@ 1.3 Differentiating the generator
A basic structural question is whether the moving generator $i=i(x,y)$, defined by
$i^{2}+\beta(x,y)i+\alpha(x,y)=0,$
can be differentiated in a meaningful way. The generator is not independent data but a chosen $C^{1}$ branch of a root of this polynomial, whose coefficients $\alpha,\beta$ lie in $C^{1}(\Omega)$. Its differentiability therefore follows by implicit differentiation of the structure relation.
Differentiating with respect to $x$ and $y$ gives
Whenever $2i+\beta$ is invertible in the fiber algebra (i.e. away from degeneracy of the quadratic), these equations uniquely determine $i_{x}$ and $i_{y}$ as algebra-valued functions depending only on $\alpha,\beta$ and their first derivatives. In the elliptic regime this invertibility is automatic, so the derivatives of the generator are explicit consequences of the structure reduction rather than additional geometric data. The computation relies only on the invertibility of the element $2i+\beta$ in the fiber algebra. This condition holds automatically in the elliptic regime and, more generally, in any nondegenerate quadratic algebra away from zero divisors. What distinguishes the elliptic case is not the algebraic solvability of these equations but the uniform positivity and invertibility properties that follow from (1.1).
@@ 1.6 Explicit Formula for the Inhomogeneous Coefficients
Using the formulas for $i_{x}$ and $i_{y}$ obtained in (1.7), we compute the intrinsic combination $i_{x}+i\,i_{y}$ in terms of $\alpha$, $\beta$ and their derivatives.
@@ 2.4 Canonical Spectral Parameter
This change of variables depends only on the structure reduction. It introduces no new assumptions and does not involve differentiation. The sign choice $b>0$ selects the branch corresponding to the elliptic orientation and ensures that $\lambda$ takes values in the upper half–plane.
@@ 3.2 The Obstruction Field
and that its derivatives $i_{x}$ and $i_{y}$ are well-defined algebra-valued sections by the implicit differentiation of the structure relation (Chapter 1).
@@ 3.4 Rigidity and Conservative Transport
$\lambda_{x}+\lambda\,\lambda_{y}=G_{0}+\lambda\,G_{1},$
where the forcing terms $(G_{0},G_{1})$ are the real coefficients of $\mathcal{G}=G_{0}+G_{1}i$ in the moving basis, as in (2.2).
@@ 3.6 The Rigidity Theorem
admits an equivalent intrinsic reformulation. It states that the generator $i(x,y)$, viewed as the generator section of the rank–two algebra bundle it defines, is holomorphic with respect to its own variable elliptic structure:
@@ 3.7 Interpretation
This reordering clarifies the role of rigidity: it is not an assumption imposed to obtain dynamics, but a compatibility condition that restores analytic integrability of the variable Cauchy–Riemann calculus within an already dynamical framework.
In chapter  5 we show that this conservative regime is precisely what allows integral representation formulas and a coherent function theory to re-emerge.
@@ 4.1 Purpose and Scope
In the previous chapters, we established that rigid variable elliptic structures are governed by a spectral transport law which, in the conservative regime, reduces to the classical complex Burgers equation.
The purpose of this chapter is to harness this well–known transport dynamics to construct concrete geometric examples. It is important to clarify the logical relationship between the analysis and the underlying PDE theory:
  • We do not aim to present new results regarding the Burgers equation itself, which is a mature and extensively studied subject in fluid dynamics and mathematical physics.
  • Instead, we adopt a constructive approach: we select elementary, standard solutions of the Burgers equation (specifically, rational profiles) and interpret them as the spectral data for a variable elliptic algebra.
By doing so, we demonstrate that rigidity is not a vacuous condition satisfied only by constants, but a rich regime capable of supporting non–trivial, moving geometries. To keep the exposition self–contained and consistent with the first–principles approach of this monograph, we briefly recall the necessary method of characteristics before deriving the explicit $\varepsilon$–family of structures.
4.2 The Burgers Substrate
In the rigid regime ($i_{x}+i\,i_{y}=0$), the universal structural transport reduces to the inviscid complex Burgers equation for the spectral parameter $\lambda$:
$\lambda_{x}+\lambda\,\lambda_{y}=0.$ (4.1)
For the purposes of this monograph, we view Equation (4.1) as a substrate—a pre–existing dynamical engine that generates the coefficients for our geometric structures. While the global theory of this equation involves shock waves and weak solutions, our geometric focus (ellipticity) restricts us to the domain of smooth solutions prior to gradient catastrophe.
4.3 Recalling the Characteristic Formulation
We utilize the standard method of characteristics to generate solutions. Equation (4.1) implies that $\lambda$ is constant along the integral curves of the vector field $(1,\lambda)$:
This classical result immediately implies the properties relevant to our construction:
  • Propagation: Structural data is transported linearly along lines in the complexified $(x,y)$ space.
  • Implicit Solutions: The general solution is given implicitly by $\lambda=F(y-\lambda x)$ for an arbitrary function $F$.
  • Blow–up: Smooth solutions generally develop singularities in finite time (characteristic crossing), which corresponds to the degeneration of the elliptic structure.
In the following sections, we apply this standard machinery to the specific ansatz of linear dependence, recovering a canonical family of rigid structures.
@@ 4.5 Characteristic Formulation
Equivalently, along integral curves of the vector field $(1,\lambda)$ in the (x,y)(x,y)–plane, the spectral parameter is conserved. This immediately implies:
@@ 4.6 Real Coefficient Representation
This system governs the evolution of the structure coefficients $(\alpha,\beta)$ in the rigid regime. It is a quasilinear first-order system of transport type.
@@ 4.8 Reduction to Ordinary Differential Equations
Substituting $a=a(x)$ and $b=b_{0}$ (constant) into (4.4), the second equation is identically satisfied, while the first reduces to
$a_{x}+a\,a_{y}=0.$
Since $a=a(x)$, this forces $a_{y}=0$, so the system imposes no further restriction on $a(x)$ at this stage. To obtain a nontrivial rigid structure, it is more transparent to work directly in terms of $(\alpha,\beta)$.
The nontrivial case is obtained by allowing $a_{x}\neq 0$ while keeping $b$ constant, which forces linear dependence in $y$.
@@
Because the rigidity condition is most naturally expressed in the original coefficients $(\alpha,\beta)$, we now translate the ansatz back into these variables.
@@ 4.10 The $\varepsilon$–Family of Rigid Structures
For every $\varepsilon\in\mathbb{R}$, the coefficients (4.7) define a rigid elliptic structure on the region where both $1-\varepsilon x\neq 0$ and the ellipticity condition (4.8) hold.
@@ 5.8 Cauchy–Pompeiu Representation Theorem
Covariant Holomorphicity and Gauge Structure in the Rigid Regime